# Local Mixtures of Experts: Essentially Free Test-Time Training via Model Merging

### v1

@@ 5 Discussion and Future Work
We propose Test-Time Model Merging (TTMM) which scales the mixture of experts paradigm to orders of magnitude more experts than in state-of-the-art multitask methods, leading to a substantial improvement in language modeling ability. We find that TTMM can achieve performance gain close to TTT. TTMM achieves this with almost the same test-time cost as inference with the base model, and without increasing train-time FLOPs or GPU memory compared to multitask fine-tuning.
In achieving close to the performance of TTT without nearly any test-time overhead, TTMM opens up several exciting directions for future research. First, TTMM makes it feasible to re-select models online during generation or evaluation based on the growing prefix, which with TTT, is too expensive in most applications. Gururangan et al. (2023) have already shown that online re-selection can work with an ensemble selected from a smaller number of experts. Second, one may further improve the performance of TTMM using tools from the literature on model merging and MoEs aimed at avoiding interference between experts and improving the routing of data to experts, e.g., by fine-tuning the router on holdout data. TTMM can further be applied to settings with privileged information by adding or removing experts dynamically based on user access. Furthermore, it may be possible to leverage the large number of experts in TTMM for uncertainty estimation, e.g., by tracking also the variance of the merged parameters. Particularly interesting would be an evaluation of TTMM on non-perplexity tasks such as Q&A or mathematics, or on a non-language domain such as image generation. Finally, many reasoning language models (called RLMs) such as DeepSeek R1 (Guo et al., 2025) are based on a MoE architecture, and TTT has been shown to be able to improve RLMs (Zuo et al., 2025). Since TTMM can be seen as a scalable MoE architecture that approximates TTT, it would be particularly interesting to study whether TTMM can improve reasoning models.
@@ Acknowledgments
{mdframed}
[ backgroundcolor=lightgray, linecolor=lightgray, linewidth=0pt, roundcorner=5pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt ] {minted}[ linenos, numbersep=4pt, xleftmargin=0.7em, frame=none, fontsize=]python for name in lora_param_names: As = torch.stack([ d[name][”A”] for d in lora_dicts]) Bs = torch.stack([ d[name][”B”] for d in lora_dicts]) delta = torch.einsum( ”k i r, k r o -¿ i o”, Bs, As)
@@ Appendix B Equivalence of Cross-Attention and RBF Kernel
We run single-step gradient descent with step size $\eta>0$ from the same initialization $\boldsymbol{\theta}$:
@@
Appendix D Evaluation of Alternative Merging Coefficients
@@
Intuitively, these weights account for redundancy between clusters. Suppose we duplicate a cluster $m$ times, then the above weights downweigh each of the $m$ duplicates by a factor $\nicefrac{{1}}{{m}}$. The cross-attention mechanism in TTMM, on the other hand, would treat each duplicate cluster as if it was a distinct cluster, and hence increase the weight of the duplicated cluster by $m$.
@@ Appendix E Experiment Details
Both the Wikipedia and the GitHub (Python) corpora contain roughly 7M documents. Therefore, since each train document is at most 1024 tokens long, each expert is trained on at most $\approx$7M tokens if we train $1000$ experts. This is a relatively small dataset size for training experts, compared to datasets used for training common MoE or multitask models.
@@ Appendix F Qualitative Examples
    Clustering: To investigate the clustering, we use an LLM to summarize each cluster’s content with a title based on 10 data points of each cluster. Tables 6 and 7 list some of those cluster titles.

### v2

@@ 5 Discussion and Future Work
We propose Test-Time Model Merging (TTMM) which scales the mixture of experts paradigm to orders of magnitude more experts than in state-of-the-art multitask methods, leading to a substantial improvement in language modeling ability. We find that TTMM can achieve performance gain close to TTT by merging few active experts into a single model per query. TTMM achieves this with almost the same test-time cost as inference with the base model, and without increasing train-time FLOPs or GPU memory compared to multitask fine-tuning. A limitation of TTMM is the additional CPU memory required for storing all experts.
In achieving close to the performance of TTT without nearly any test-time overhead, TTMM opens up several exciting directions for future research. First, TTMM makes it feasible to re-select models online during generation or evaluation based on the growing prefix, which with TTT, is too expensive in most applications. Gururangan et al. (2023) have already shown that online re-selection can work with an ensemble selected from a smaller number of experts. Second, one may further improve the performance of TTMM using tools from the literature on model merging and MoEs aimed at avoiding interference between experts and improving the routing of data to experts, e.g., by fine-tuning the router on holdout data. TTMM can further be applied to settings with privileged information by adding or removing experts dynamically based on user access. Furthermore, it may be possible to leverage the large number of experts in TTMM for uncertainty estimation, e.g., by tracking also the variance of the merged parameters. Finally, many reasoning language models (called RLMs) such as DeepSeek R1 (Guo et al., 2025) are based on a MoE architecture, and TTT has been shown to be able to improve RLMs (Zuo et al., 2025). Since TTMM can be considered a scalable MoE architecture that approximates TTT, it would be particularly interesting to study whether TTMM can improve reasoning models.
@@ Acknowledgments
for name in lora_param_names: As = torch.stack([ d[name][”A”] for d in lora_dicts]) Bs = torch.stack([ d[name][”B”] for d in lora_dicts]) delta = torch.einsum( k i r, k r o -> i o”, Bs, As)
@@ Appendix B Equivalence of Cross-Attention and RBF Kernel
We first analyze single-step gradient descent with step size $\eta>0$ from the same initialization $\boldsymbol{\theta}$:
@@
To complete the proof of 3.1, we simply observe that the bound $\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(1)}-\boldsymbol{\theta}^{\prime(1)}\|\leq\eta G(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime}))$ can be applied recursively from the shared initialization $\boldsymbol{\theta}^{(0)}$ for $T$ steps of gradient descent via the relation
$\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(t+1)}-\boldsymbol{\theta}^{\prime(t+1)}\|\leq\|\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}^{(t)}-\boldsymbol{\theta}^{\prime(t)}\|+\eta G(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$
Unrolling this recursion proves 3.1 with an additional factor $T$ compared to Proposition C.1.
Appendix D Ablations
D.1 Fixing the Number of Active Experts
In Table 3, we ablate the choice of selecting the number of active experts dynamically per prompt depending on the threshold $\tau$ (cf. Algorithm 2) rather than specifying a fixed sparsity $N$. On our evaluated corpora, we find no significant difference in performance.
Model # Active Experts Wikipedia GitHub (Python)
TTMM Fixed TTMM Fixed
Llama-3.2-1B 1 7.669 7.669 2.552 2.552
3 7.571 7.580 2.519 2.505
10 7.510 7.509 2.492 2.491
Qwen2.5-1.5B 1 7.166 7.166 2.330 2.331
3 7.101 7.104 2.304 2.295
10 7.080 7.080 2.287 2.287
Table 3: Perplexity comparison (lower is better) between TTMM (dynamic expert routing) and fixed expert selection for Llama and Qwen across different numbers of active experts.
D.2 Different Embedding Models
In Table 4, we evaluate different embedding models on MMLU and find that TTMM outperforms the base model regardless of the choice of embedding model. Notably, the largest Qwen2-1.5B embedding model performs worse than the 100M MPNet model. A possible reason is that finding a decent clustering might be significantly easier than directly solving questions, since the former only requires a broad understanding of the meaning of some keywords while the latter requires some ability of reasoning. Stella (Zhang et al., 2024a) performs the best among the tested embedding models.
Model Humanities Social Sciences STEM Other Overall
Fine-tuned 44.80 54.24 40.79 54.43 48.10
all-mpnet-base-v2 45.95 54.76 40.56 54.72 48.61
stella_en_400M_v5 45.46 55.31 41.77 55.26 48.96
gte-Qwen2-1.5B-instruct 45.27 54.76 40.91 54.65 48.45
Table 4: MMLU accuracy (in %) comparing TTMM with different embedding models against the fine-tuned baseline. Bold indicates the highest accuracy.
D.3 Evaluation of Alternative Merging Coefficients
@@
Intuitively, these weights account for redundancy between clusters. Suppose we duplicate a cluster $m$ times, then the above weights downweigh each of the $m$ duplicates by a factor $\nicefrac{{1}}{{m}}$. The cross-attention mechanism in TTMM, on the other hand, would treat each duplicate cluster as if it were a distinct cluster, and hence increase the weight of the duplicated cluster by $m$.
@@ Appendix E Experiment Details
Both the Wikipedia and the GitHub (Python) corpora contain roughly 7M documents. Therefore, since each train document is at most 1024 tokens long, each expert is trained on at most $\approx$7M tokens if we train 100010001000 experts. This is a relatively small dataset size for training experts, compared to datasets used for training common MoE or multitask models.
@@ Appendix F Qualitative Examples
    Clustering: To investigate the clustering, we use an LLM to summarize each cluster’s content with a title based on 10 data points of each cluster. Tables 9 and 10 list some of those cluster titles.